# Investigating Coverage Criteria in Large Language Models: An In-Depth Study Through Jailbreak Attacks

### v1

@@ V-A Effectiveness Comparison
Training Set 93.92% 96.28% 100.00%
Testing set 94.90% 96.10% 98.00%
TruthfulQA 94.12% 91.19% 87.03%
Adversarial Attack 92.00% 84.50% 82.50%
TABLE V: Experimental Results of Perplexity Filter and PARDEN
DataSet Perplexity PARDEN
OPT-125M
Llama-2-7B-Chat
Pythia-12B
OPT-125M
Llama-2-7B-Chat
Pythia-12B
Sentence Window Sentence Window Sentence Window Sentence Window Sentence Window Sentence Window
Alpaca-gpt4 98.63% 100.00% 98.70% 97.83% 97.20% 100.00% 83.57% 93.40% 57.37% 54.70% 85.20% 85.00%
JailBreakV 8.97% 0.07% 9.43% 10.03% 9.40% 9.57% 46.13% 37.63% 54.60% 61.90% 6.00% 3.53%
TruthfulQA 99.76% 100.00% 97.55% 99.39% 98.78% 100.00% 90.58% 96.33% 38.92% 34.52% 83.60% 81.64%
Adversarial Attack 67.00% 5.00% 81.50% 78.00% 80.00% 59.50% 31.00% 27.50% 86.50% 89.50% 17.50% 17.00%
Evaluation results: Experimental results are presented in Tables IV and V. Table IV demonstrates that the classifier based on NC performs well across three models. Specifically, the classifier achieves an average classification accuracy of 96.33% on the testing set, indicating excellent performance in distinguishing between normal and attack queries after training. Furthermore, on the TruthfulQA dataset, the classifier achieves an average classification accuracy of 90.78% across the three models. The classifier’s effectiveness on the TruthfulQA dataset, which includes questions prone to errors, demonstrates its robustness and generalization ability to normal queries. For adversarial attack queries, the classifier achieves an average accuracy of 86.33%, further showcasing its generalization capability. This is particularly notable considering the different attack patterns in the datasets: GCG involves gradient-based adversarial attack suffixes, while JailBreakV comprises template-based, persuasion-related jailbreak attack queries.
A further comparison with two baselines in Table V reveals that the perplexity filter performs well in detecting normal queries, achieving nearly 100% accuracy with a window size of 10. However, its performance declines significantly on JailBreakV and adversarial attack queries, with an average classification accuracy of only 7.91% on JailBreakV. For adversarial attack queries, the perplexity filter achieves sentence-level classification accuracies of 81.5% on the Llama-2-7B-Chat model, 80% on the Pythia-12B model, and only 67% on the OPT-125M model. This indicates that the perplexity filter struggles with low-complexity attack queries, such as template and instruction-based jailbreak attacks in JailBreakV.
PARDEN’s performance varies across datasets and models. It effectively detects normal queries on OPT-125M and Pythia-12B, achieving 93.4% window-based accuracy on Alpaca-gpt4 for OPT-125M and 85.2% sentence-level accuracy on Pythia-12B, with similar results on TruthfulQA. However, both models perform poorly on attack queries, possibly due to insufficient safety alignment, as OPT-125M includes unfiltered internet content, and Pythia-12B is designed for interpretability research, not practical deployment. On Llama-2-7B-Chat, PARDEN only performs well on adversarial attack queries, with 89.5% sentence-level accuracy.
Besides superior detection performance, our method exhibits three advantages: ❶ The detection method is lightweight as it utilizes only a small set of features, i.e. the number of activated neurons in each layer, as the input features. ❷ Our method is independent of query length, enabling it to flexibly address complex and variable attack methods. In contrast, the detailed hyper-parameters and performance of previous methods, such as the perplexity-based method, vary according to the length of the query being detected. ❸ Compared to other output detection methods, our approach completes detection as soon as the model generates the first token, avoiding the resource waste associated with waiting for the full output. This key feature can significantly save resources and computations compared to methods that only identify jailbreak attacks after the complete generation.
Findings: The NC-based classifier surpasses the perplexity filter and PARDEN in detecting attack queries, showing strong generalization and adaptability independent of model architecture or alignment. Moreover, it identifies attack queries before output generation, saving computational resources.
@@ VI Threats to Validity
External Threats: External validity threats arise from the specific settings chosen in our study, which may raise concerns about the generalizability of our proposed method. To mitigate these threats, we select a variety of evaluation settings. Specifically, we use four large language models with different architectures and parameters: OPT-125M, Llama-2-7B-Chat, Pythia-12B, and Gemma-2-27B-it. Additionally, we employ four datasets—Alpaca-gpt4, JailBreakV-28k, TruthfulQA, and AdvBench—that cover a wide range of test scenarios. Our study is based on five widely used coverage criteria: NC, TKNC, TKNP, TFC, and NLC. These choices aim to enhance the generalizability of our findings to common scenarios in this field.
@@
Internal Threats: Internal validity threats stem from the tools used in our study, including llm-attacks (the official implementation of GCG). Moreover, accurately reproducing the coverage criteria and the perplexity filter and PARDEN for input detection also presents potential threats. Inherent randomness in model training further poses a potential threat to internal validity. We address this by repeating key experiments more than three times to report the average results.
VII Related Work
LLM Testing and Evaluation. Previous studies have proposed several benchmarks to comprehensively evaluate the fundamental capabilities of LLMs in a standardized manner. For example, MMLU [29] aims to measure the multitask accuracy of text models across 57 tasks, including mathematics, history, computer science, and law. Big-Bench [30] assesses language models across 204 tasks, covering areas such as linguistics, child development, mathematics, commonsense reasoning, biology, physics, social bias, and software development. LMentry [31] evaluates the capabilities and robustness of LLMs through a set of simple tasks for humans, including sentence construction, part-of-speech recognition, and word length comparison. AGIEval [32] assesses model proficiency through standardized human exams, such as college entrance exams and math competitions. These benchmarks collectively establish a foundation for evaluating and enhancing LLM reliability and performance across various domains and tasks.
LLM Defenses. Defensive mechanisms for LLMs are also evolving. For example, erase-and-check [33] is a certified safe defense framework against adversarial prompts, which defends by sequentially removing tokens from the prompt and using a safety filter to inspect the resulting subsequences. SmoothLLM [34] and JailGuard [35] defend by randomly perturbing multiple copies of an input prompt and aggregating the corresponding predictions, significantly reducing the success rate of jailbreak attacks. Other defense methods include perplexity-based detection [27, 36] and self-classification by the LLM [28, 37], which detect disordered characters in prompts or classify input prompt harmfulness directly through the model.
VIII conclusion
@@ VII Related Work
In this study, we conducted an extensive empirical investigation into the effectiveness of traditional coverage criteria in LLMs across three key dimensions: criterion level, layer level, and token level. Our findings reveal significant differences in neuron coverage between normal and malicious queries, highlighting the potential of these criteria in identifying abnormalities in LLMs. Building upon these insights, we developed a novel downstream application for real-time detection of malicious attacks based on neural activation features, achieving an impressive accuracy of 96.33%. Our findings enhance the understanding of security testing in LLMs and lay a methodological foundation for developing robust AI applications using neural activation features to detect malicious attacks.
Broader Impact. Our paper demonstrates the effectiveness of coverage criteria on LLMs, focusing on jailbreak attacks due to their widely recognized impact and well-defined nature. We hope to inspire the community to explore coverage criteria effectiveness across a broader range of security and functional issues in future research.

### v3

@@ V-B Test Case Prioritization
Synonymous 91.60% 83.40% 72.60%
JailbreakV 77.60% 97.80% 94.60%
Avg 84.60% 90.60% 83.60%
Test case prioritization represents another critical application. By leveraging coverage levels, test cases likely to reveal model faults (high coverage) are prioritized, while redundant ones (low coverage) are filtered, improving testing efficiency and reducing resource usage.
Setup: We extend the experiments conducted in RQ1 by defining the average coverage increase of individual synonymous queries (redundant test cases) relative to $S_{N}$ as a threshold. Test cases with coverage increases below this threshold are considered redundant, whereas those exceeding the threshold are identified as prioritized test cases. To evaluate the effectiveness of this method, we conduct experiments using 500 synonymous queries (redundant test cases) and 500 attack queries (prioritized test cases, extracted from JailBreakV).
Evaluation results: Table VI shows the classification accuracies of this method on three LLMs. For redundant test cases (synonymous queries), the method achieves high accuracies of 91.60% on OPT-125M, 83.40% on Llama-2-7B-Chat, and 72.60% on Pythia-12B, effectively filtering them out. For prioritized test cases (attack queries), it attains accuracies of 77.60%, 97.80%, and 94.60%, respectively, demonstrating strong capability in identifying test cases likely to uncover model faults. The overall average accuracies—84.60% for OPT-125M, 90.60% for Llama-2-7B-Chat, and 83.60% for Pythia-12B—validate the effectiveness of coverage criteria in prioritizing test cases, enabling efficient vulnerability detection while reducing redundancy.
V-C Jailbreak Case Generation
Figure 6: Comparison of Successful Jailbreak Queries Generated by Coverage-Guided and Random Methods Over Iterations.
By utilizing coverage to guide the creation of attack examples, Jailbreak Case Generation method identifies areas of the model that remain unexplored. Iterative refinement of prompts based on coverage gains ensures that the generated cases are effective in exposing vulnerabilities while promoting diversity among test cases. Here, we conduct a preliminary exploration to showcase the potential of coverage-guided jailbreak case generation.
Setup: We use Llama-2-7b-chat as the target model, initializing with five jailbreak queries as seeds. Over five iterations, GPT-4 generates ten new jailbreak queries per round through prompt rewriting. The query with the highest coverage increase is selected as the next seed in the coverage-guided approach. For comparison, a random strategy selects seeds randomly from rewritten candidates. Each method ultimately generates 250 new jailbreak queries to evaluate effectiveness.
Evaluation results: As shown in Figure 6, the coverage-guided method outperforms the random strategy in generating successful jailbreak queries. In Round 1, our approach generates 6 successful queries compared to 2 by the random method. This gap widens in subsequent rounds, with our method producing 11, 19, 24, and 32 successful queries in Rounds 2 to 5, respectively, while the random strategy yields 2, 6, 9, and 12. These results confirm that coverage gains effectively guide query generation toward unexplored areas, increasing the number of successful jailbreak cases and enhancing robustness testing for LLMs.
@@ VI Threats to Validity
External Threats: External validity threats arise from the specific settings chosen in our study, which may raise concerns about the generalizability of our proposed method. To mitigate these threats, we select a variety of evaluation settings. Specifically, we use four large language models with different architectures and parameters: OPT-125M, Llama-2-7B-Chat, Pythia-12B, and Gemma-2-27B-it. Additionally, we employ four datasets—Alpaca-gpt4, JailBreakV-28k, TruthfulQA, and AdvBench—that cover a wide range of test scenarios. Our study is based on five widely used coverage criteria: NC, TKNC, TKNP, TFC, and NLC. The choices aim to enhance the generalizability of our findings in this field.
@@
Internal Threats: Internal validity threats stem from the tools used in our study, including llm-attacks (GCG), DeepInception, Masterkey, and Self-Reminder. Moreover, accurately reproducing the coverage criteria and the perplexity filter and PARDEN for input detection also presents potential threats. Inherent randomness in model training further poses a potential threat to internal validity. We address this by repeating key experiments more than three times to report the average results.
VII conclusion
@@
In this study, we conduct an extensive empirical investigation into the effectiveness of traditional coverage criteria in LLMs across three key dimensions: criterion level, layer level, and token level. Our findings reveal significant differences in neuron coverage between normal and malicious queries, highlighting the potential of these criteria in identifying abnormalities in LLMs. Building upon these insights, we explore three downstream applications based on coverage criteria: Real-time Jailbreak Detection, Test Case Prioritization, and Jailbreak Case Generation, all of which achieve outstanding performance. Our findings enhance the understanding of security testing in LLMs, highlight the potential of coverage criteria in addressing broader security and functional challenges, and lay a methodological foundation for developing robust AI applications using neural activation features.
Acknowledgments
We thank the anonymous reviewers for their constructive comments. This work was supported by the National NSF of China (grants No.62302176, No.62072046, 62302181), the Key R&D Program of Hubei Province (2023BAB017, 2023BAB079), the Knowledge Innovation Program of Wuhan-Basic Research (2022010801010083), and the National Research Foundation, Prime Minister’s Office, Singapore under the Campus for Research Excellence and Technological Enterprise (CREATE) programme.